OpenStudy (anonymous):

What is the vertex and the equation of the axis of symmetry of the graph of y=x^2-6x-7?

OpenStudy (ashes.boom):

okay so to find the x point of the vertex you use the equation x = -b/2a using the coefficients from the equation: a = 1, b = -6 and c = -7 so your equation would look like this: x = -(-6)/2(1) then you just solve it x = 6/2 x = 3 now that you have the x value you can put it in the equation for y

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (welshfella):

another way to do this is by rewriting the equation in vertex form x^2 - 6x - 7 = (x - 3)^2 - 9 - 7 = (x - 3)^2 - 16 this shows that the vertex is at (3,-16) and axis of symmetry is x = 3